OpenStudy (anonymous):

Write the equation in slope intercept form: Through (-5, 4); slope 1/2.

OpenStudy (anonymous):

y = mx+c slope is m and you can find value of c by putting (-5,4 ) in the equation

myininaya (myininaya):

Equation of line can be written in y=mx+b form where m is the slope and b is the y-intercept you are given the slope so go ahead and replace the m here with 1/2 You know a point on the line so replace (x,y) with (-5,4) and solve for b.

OpenStudy (anonymous):

I put: 4=1/2 (-5) + b (just plugging in the given points) and then, I multiply the -5 by the 1 on 1/2? Which makes it -5/2, correct?

OpenStudy (ranga):

You have 4 = -5/2 + b. Multiply both sides by 2 and solve for b.

OpenStudy (anonymous):

So when I have a fraction as the slope, I just multiply it by the denominator along with whatever the y is at all times?

OpenStudy (ranga):

Multiplying by the denominator helps get rid of the fraction part and makes it easy to solve the equation for b.

OpenStudy (anonymous):

How come the answer key says that the equation ends up being: y=1/2x + 13/2? I see how to get the 13 by multiplying the 2 with the -5/2 and the 4, but how does 13 become 13/2?

OpenStudy (ranga):

4 = -5/2 + b Multiply each term by 2. 8 = -5 + 2b 2b = 8 + 5 = 13 b = 13/2 Slope intercept form is y = mx + b m = slope = 1/2 b = 13/2 So y = 1/2x + 13/2

OpenStudy (anonymous):

Thank you so much! I understand where the 13 turns into 13/2 now.

OpenStudy (ranga):

You are welcome!